OpenStudy (anonymous):

Sally applies a total force of 299 N with a rope to drag a wooden crate of mass 100 kg across a floor with a constant acceleration in the x-direction of a = 0.1 m/s 2 . The rope tied to the crate is pulled at an angle of θ = 59 ◦ relative to the floor. m Calculate the total work done by friction as the block moves a distance 21.8 m over the horizontal surface. The acceleration due to gravity is 9.8 m/s 2 . Answer in units of J

OpenStudy (anonymous):

net force = ma let the frictional force be denoted by R therefore, Applied force - frictional force = mass x acceleration 299cos59 - R = 100 x 0.1 hence, 153.9963844 - R = 10 -R = -143.9963844 thereforce R (frictional force) = 143.9963844 workdone by friction = force x distance 143.9963844N x 21.8m =3139.12118J